OpenStudy (sissyedgar):

Solve the equation. x² = 150 A. ±25√6 B. 15 C. -15, 15 D. ±5√6 I know It's not B. or C.

OpenStudy (sissyedgar):

Lol ,i was just about to tag you

OpenStudy (abdullahm):

What are the factors of 150?

rebeccaxhawaii (rebeccaxhawaii):

B and C are incorrect because 15*15 does not equal 150.

rebeccaxhawaii (rebeccaxhawaii):

Put two factors of 150. Make sure that one can be squared evenly.

OpenStudy (sissyedgar):

1 - 2 - 3 - 5 - 6 - 10 - 15 - 25 - 30 - 50 - 75 - 150

OpenStudy (abdullahm):

Here, let me do an example for you: \(\Large\bf \sqrt{250} = \sqrt{2 \times 5\times 5\times 5} = \sqrt{5^2 \times 2 \times 5} =\\\Large\bf 5\sqrt{2\times 5} = 5\sqrt{10}\) When you have something square under the square root, you bring it out but without the square

rebeccaxhawaii (rebeccaxhawaii):

Here's something a little less overwhelming \[x^2=250\] \[250=25*10\] you are able to square 25 perfectly \[5\sqrt{10} \]

OpenStudy (sissyedgar):

Wait.. So Is it D?

OpenStudy (abdullahm):

Correct!

rebeccaxhawaii (rebeccaxhawaii):

Explain to me why it would be D. Not saying if it right or wrong.

rebeccaxhawaii (rebeccaxhawaii):

@AbdullahM let him show his work

rebeccaxhawaii (rebeccaxhawaii):

I wanna see if he understands.

OpenStudy (sissyedgar):

I got D becuase.... I did (√6 x 5)² And.. Im a girl.. J.s.

rebeccaxhawaii (rebeccaxhawaii):

My apologies. I saw "sissy" and realized you are a girl. AND GOOD JOB!!!!!

OpenStudy (sissyedgar):

lol, Its okay! And thankz!